How do I solve this quadratic equation?It is not a textbook question. It is from a handmade study guide. I would appreciate the help!!! Thanks.
2x^2-4x=0

Answer by nerdybill(7384)

2x^2-4x=0
.
Begin by factoring the left.
Do this by terming the largest common factor -- happens to be 2x:
(2x)(x-2)=0
.
Now, if either (2x) OR (x-2) equals zero, the entire left side would be zero making the statement true.
.
Therefore, your two solutions are found by setting both terms to zero and solving for 'x':
solution 1:
2x = 0
x = 0
.
solution 2:
x-2 = 0
x = 2
.
Solution set:
x = {0, 2}